If the half life of radio active is T then the fraction that would remain after a time T/2 is

Grade:12th pass

1 Answers

Yuvraj Singh
2 years ago
If T is the half-life of a radioactive material, then the fraction that would remain after a time (T)/(2) is. ∴NN0=(12)T/2T=(12)1/2=1√2